Use the given conditions to write an equation for the line in point-slope form.
Slope=4 passing through (-3,7)
Slope=6/7 passing through (8,7)



Answer by stanbon(75887)

Use the given conditions to write an equation for the line in point-slope form.
Slope=4 passing through (-3,7)
Form: y = mx+b
You know m = 4 and y = 7 when x = -3 ; solve for "b":
7 = 4*-3 + b
b = 14
Equation:
y = 4x + 14
===================
Slope=6/7 passing through (8,7)
Form: Y = mx+b
7 = (6/7)8 + b
7 = (48/7) + b
49/7 = 48/7 + b
b = 1/7
-----
Equation:
y = (6/7)x + (1/7)
==================